# Barbie: Text to Barbie-Style 3D Avatars

### v6

@@ Abstract
To integrate digital humans into everyday life, there is a strong demand for generating high-quality, fine-grained disentangled 3D avatars that support expressive animation and simulation capabilities, ideally from low-cost textual inputs. Although text-driven 3D avatar generation has made significant progress by leveraging 2D generative priors, existing methods still struggle to fulfill all these requirements simultaneously. To address this challenge, we propose Barbie, a novel text-driven framework for generating animatable 3D avatars with separable shoes, accessories, and simulation-ready garments, truly capturing the iconic “Barbie doll” aesthetic. The core of our framework lies in an expressive 3D representation combined with appropriate modeling constraints. Unlike previous methods, we innovatively employ G-Shell to uniformly model both watertight components (e.g., bodies, shoes, and accessories) and non-watertight garments compatible with simulation. Furthermore, we introduce a well-designed initialization and a hole regularization loss to ensure clean open surface modeling. These disentangled 3D representations are then optimized by specialized expert diffusion models tailored to each domain, ensuring high-fidelity outputs. To mitigate geometric artifacts and texture conflicts when combining different expert models, we further propose several effective geometric losses and strategies. Extensive experiments demonstrate that Barbie outperforms existing methods in both dressed human and outfit generation. Our framework further enables diverse applications, including apparel combination, editing, expressive animation, and physical simulation. Our project page is: https://xiaokunsun.github.io/Barbie.github.io
Index Terms:
3D disentangled avatar generation, expressive animation, physical simulation, diffusion model, score distillation
{strip}[Uncaptioned image]
Figure 1: Our method generates Barbie-style 3D avatars from textual input. “Barbie-style” refers to the following key characteristics: (1) High-Quality geometry and realistic appearance, ensuring visually lifelike avatars; (2) Fine-Grained Decoupling, separating body, clothing, shoes, and accessories to enable flexible apparel combination and editing; (3) Expressive Animation, supporting a wide range of body movements, facial expressions, and hand gestures; (4) Simulation Compatibility, enabling modeling of non-watertight garments and seamless integration into existing physical simulation pipelines.
@@ I Introduction
It is natural to ask: Can we automatically generate Barbie-like 3D avatars based on text? Our answer is yes, but this is a non-trivial task due to several challenges. (1) How do we model watertight components (e.g., bodies, shoes, and accessories) and simulation-ready garments that are non-watertight? While more expressive representations (e.g., G-Shell [63]) offer potential solutions, initializing and regularizing them to accurately represent open surfaces without multi-view image inputs remains an open problem. (2) Providing suitable constraints for optimizing decoupled bodies and outfits to achieve domain-specific realism is both key and challenging.
@@
To address these challenges, we introduce Barbie, a novel text-driven framework for generating Barbie-style 3D avatars. As shown in Fig. 1, avatars generated by Barbie not only exhibit exquisite geometry and textures but also feature a variety of decoupled shoes, accessories, and simulation-ready garments that can be freely combined and edited. Furthermore, our framework supports strong animation expressiveness and compatibility with physical simulations, thereby satisfying all four aforementioned requirements simultaneously.
The core of the framework consists of two key components: (1) Unlike previous methods, we employ expressive G-Shell [63] to uniformly model both watertight components (e.g., bodies, and accessories) and non-watertight garments. Additionally, we design an efficient initialization strategy and a hole-preserving loss to achieve clean and well-defined open surface modeling. (2) Instead of relying on a single general model as in prior text-to-disentangled-avatar methods, we appropriately incorporate different expert diffusion models to guarantee the domain-specific fidelity. A series of effective regularization losses and strategies is also proposed to address geometric artifacts and texture conflicts when combining different expert models. Specifically, our method generates Barbie-like 3D digital humans in three stages: First, we generate a reasonable and realistic base body using human-specific diffusion models along with a proposed SMPLX-evolving prior loss. Second, we initialize apparel with the semantic-aligned body and optimize it using object-specific generative priors and several solid geometric losses. Finally, we jointly fine-tune the assembled avatar to enhance texture harmony and consistency. Through this pipeline, the generated 3D avatars are animatable and disentangled, featuring high-quality bodies, shoes, accessories, and simulation-ready garments, delivering a truly immersive “Barbie-style” digital experience.
@@
  • We introduce Barbie, a novel text-driven framework for generating realistic and highly disentangled 3D avatars. The framework enables the decoupling of bodies, garments, shoes, and accessories, while enabling outfit transfer, editing, expressive animation, and simulation.
@@
  • Extensive experiments demonstrate that Barbie outperforms existing methods in avatar and outfit generation, achieving superior results in geometry quality, texture detail, and alignment with text descriptions.
@@ II Related Work
Text to Holistic 3D Avatar Generation. The field of text-to-3D generation has witnessed significant progress [12, 14, 19, 20, 15, 21, 18, 16], largely driven by advances in text-to-image models [10, 9, 62]. Building upon the success in generating general 3D objects, various methods [22, 24, 31, 30, 25, 32, 26, 64, 23, 36, 80, 34, 27] have been proposed to generate complex 3D avatars. AvatarCLIP [22] pioneers zero-shot generation of 3D digital humans from text prompts by leveraging human priors and CLIP [2]. With the introduction of the Score Distillation Sampling (SDS) loss [12], several subsequent works [24, 31, 30, 25] combine it with parametric human models [41, 42, 61] to significantly improve the fidelity of generated avatars. To enhance animation expressiveness and simulation compatibility, approaches [32, 26, 64, 23, 36, 80] replace implicit NeRF [47, 48] with explicit representations such as 3DGS [50] and SMPL-X [42] for modeling 3D humans. However, purely implicit or explicit representations struggle to capture complex geometric details, limiting the quality of generated digital humans. Accordingly, HumanNorm [34] and SeeAvatar [27] propose using the hybrid representation DMTet [51], which combines the advantages of both implicit and explicit representations to achieve high-fidelity avatar generation. Nevertheless, these methods typically model the human body, garments, shoes, and accessories as a single model, sacrificing flexibility and controllability for downstream applications such as apparel composition and physical simulation.
@@
In contrast, the avatars generated by Barbie not only exhibit exquisite geometry and realistic appearance but also wear multiple separable, realistic outfits, while supporting expressive animation and physical simulation. We summarize the main differences between Barbie and existing methods in Table I.
@@ III-B Overview
Given a text prompt, Barbie aims to create an animatable, disentangled 3D avatar dressed in simulation-ready garments, along with diverse shoes and accessories, resembling iconic Barbie dolls. Specifically, our framework consists of three stages: (1) Human Body Generation: This stage generates a reasonable and realistic basic human body by leveraging human-specific generative priors and a novel SMPLX-evolving prior loss (Sec. III-C and Fig. 2); (2) Apparel Generation: This stage models high-quality garments, shoes, and accessories piece by piece, utilizing object-specific diffusion models together with several initialization strategies and geometric losses (Sec. III-D and Fig. 3); and (3) Unified Texture Refinement: This stage enhances visual harmony and consistency by jointly fine-tuning the composed avatar (Sec. III-E and Fig. 3).
@@ III-C Human Body Generation
Since the human body is a closed surface, we utilize G-Shell ${\theta}_{h}$ in watertight mesh mode (i.e., ignoring mSDF and retaining only SDF) and a texture field ${\psi}_{h}$ to model the geometry and appearance of the human body, respectively. These components are optimized under human-specific generative priors and a novel SMPLX-evolving prior loss, as illustrated in 2.
Human Body Geometry Modeling. To enhance the robustness of 3D human generation, we start by initializing the human G-Shell ${\theta}_{h}$ with an SMPL-X model $M_{init}$. This SMPL-X mesh $M_{init}$ will be used in subsequent stages to provide rich human priors and ensure a semantic-aligned representation for outfit creation. As discussed in the introduction, using a general T2I model struggles to provide domain-specific constraints necessary for creating realistic human bodies or outfits. To address this limitation, we employ human-specific diffusion models from HumanNorm [34], which is fine-tuned with high-fidelity human data.
@@
To address this problem, we propose the SMPLX-evolving prior loss inspired by the evolving constraint introduced in SeeAvatar [27]. Specifically, we freeze the SMPL-X shape parameters $\beta$ but enhance $M_{init}$ to an evolvable $\hat{M}_{init}$ by adding learnable vertex-wise offsets. These offsets are periodically fitted to the current body every $\delta$ iteration, allowing the optimized offsets to model complex body features. The fitting objective is as follows:
@@
where $\mathcal{L}_{chamf}$ is the Chamfer distance between $\hat{M}_{init}$ and the current body shape, $\mathcal{L}_{edge}$, $\mathcal{L}_{nor}$ and $\mathcal{L}_{lap}$ are the edge length regularization, normal consistency loss, and Laplacian smoothness, respectively. The advantages of this loss are twofold: (1) Enhanced Modeling Capabilities: Compared to the smooth $M_{init}$, our $\hat{M}_{init}$ excels at capturing detailed geometric features, such as muscle contours (see the zoomed area at the bottom of Fig. 2), and offers reliable yet diverse priors for subsequent geometry generation (Fig. 8-(a)). (2) Topology Preservation: Unlike the evolving templates with arbitrary topology in SeeAvatar [27], our $\hat{M}_{init}$ preserves the SMPL-X model’s topology and semantics during the evolution process. This makes it particularly suitable for apparel initialization, composition, animation, and simulation (Fig. 9.
@@ III-D Apparel Generation
Given an SMPLX-aligned basic body generated in the previous stage, we proceed to dress the Barbie avatar in various outfits. Similar to body modeling, we employ G-Shell ${\theta}_{a}$ for apparel geometry and a texture field ${\psi}_{a}$ for apparel texture. Closed shoes and accessories are represented using G-Shell in watertight mesh mode (i.e., ignoring mSDF and retaining only SDF), while open garments are modeled using G-Shell in non-watertight mesh mode (i.e., retaining both SDF and mSDF). As illustrated in Fig. 3, the apparel generation process begins by initializing apparel using predefined SMPL-X masks that cover over a dozen daily outfits. We then create high-quality apparel piece by piece using object-specific generative priors and well-designed geometric losses. Lastly, we fine-tune the assembled avatar to improve appearance harmony and consistency.
@@
Apparel Initialization. Due to the lack of reliable multi-view image input [63, 65], initialization plays a critical role in text-driven G-Shell optimization. Therefore, we carefully design two strategies for initializing open clothes and closed shoes/accessories, enabling accurate topology modeling for each category. (1) Closed Surface Initialization: We crop the open sub-mesh of the human body using the corresponding SMPL-X mask, expand it along vertex normals to create inner and outer layers, and sew them to form a double-layer closed template mesh $M_{temp}$. This mesh initializes the SDF $s_{\theta_{a}}(\cdot)$ of the apparel G-Shell ${\theta}_{a}$. Fig. 4-(a) shows this process. (2) Open Surface Initialization: Since mSDF is defined on a watertight surface, we first initialize the SDF $s_{\theta_{a}}(\cdot)$ of ${\theta}_{a}$. Concretely, we expand the cropped open mesh, fill its holes, and obtain a single-layer watertight template mesh $M_{temp}$. Unlike the double-layer template mesh in closed surface initialization, this simpler template mesh is better suited for mSDF initialization. The initialization of mSDF is the core of the open surface initialization process. A straightforward way is to initialize mSDF according to its definition (i.e., the signed geodesic distance to the open surface boundary on the watertight template). Nonetheless, this geodesic-based approach is not only computationally expensive (5 hours) but also unstable, often causing unsmooth boundaries and floating triangles (Fig. 4-(c)). This is because geodesics are defined on the surface and cannot stably supervise a 3D implicit field.
To solve these issues, we propose a novel SDF-based mSDF initialization strategy. Specifically, we fit a pie-shaped watertight mesh $M_{pie}$ to each hole that covers its boundaries (see the Supp. Mat. for fitting details). We then use $M_{pie}$’s SDF values to initialize the mSDF $\hat{s}_{\theta_{a}}(\cdot)$ of ${\theta}_{a}$. In this way, we can remove intersections between $M_{temp}$ and $M_{pie}$, thereby obtaining the desired open surface (see the black dotted box in Fig. 4-(b)). Fig. 4-(b) illustrates the whole process. Our method offers three key advantages: (1) Efficiency: Computing SDF takes only about 3 minutes, significantly faster than computing geodesics ( 5 hours). (2) Stability: The SDF defined in 3D space can stably supervise the 3D implicit field, thus effectively avoiding floating triangles and unsmooth boundaries. (3) Feature Mesh: $M_{pie}$ provides regularization constraints for subsequent optimization (see Eq. 14). Fig. 4-(c) compares the two initialization methods.
@@
where $y_{a}$ is the description of a single outfit, and $n^{a}$ and $d^{a}$ are the rendered normal and depth maps of the apparel in the undressed state. Besides, $\mathcal{L}^{{sdn}}_{SDS}$ is the vanilla Stable Diffusion (SD) SDS loss enforced on the rendered apparel normal maps. As shown in RichDreamer [18], naive SD helps object-specific diffusion models produce more stable results. In this way, the object-specific diffusion models provide powerful guidance to modeling high-fidelity outfits (Fig. 8-(b)).

### v7

# DreamBarbie: Text to Barbie-Style 3D Avatars

@@ Abstract
To integrate digital humans into everyday life, there is a strong demand for generating high-quality, fine-grained disentangled 3D avatars that support expressive animation and simulation capabilities, ideally from low-cost textual inputs. Although text-driven 3D avatar generation has made significant progress by leveraging 2D generative priors, existing methods still struggle to fulfill all these requirements simultaneously. To address this challenge, we propose DreamBarbie, a novel text-driven framework for generating animatable 3D avatars with separable shoes, accessories, and simulation-ready garments, truly capturing the iconic “Barbie doll” aesthetic. The core of our framework lies in an expressive 3D representation combined with appropriate modeling constraints. Unlike prior methods, we use G-Shell to uniformly model watertight components (e.g., bodies, shoes) and non-watertight garments. By reformulating boundaries as Euclidean field intersections instead of manifold geodesics, we propose an SDF-based initialization and a hole regularization loss that together achieve a $100\times$ speedup and stable open topology without image input. These disentangled 3D representations are then optimized by specialized expert diffusion models tailored to each domain, ensuring high-fidelity outputs. To mitigate geometric artifacts and texture conflicts when combining different expert models, we further propose several effective geometric losses and strategies. Extensive experiments demonstrate that DreamBarbie outperforms existing methods in both dressed human and outfit generation. Our framework further enables diverse applications, including apparel combination, editing, expressive animation, and physical simulation. Project page: https://xiaokunsun.github.io/DreamBarbie.github.io/.
@@ I Introduction
It is natural to ask: Can we automatically generate Barbie-like 3D avatars based on text? Our answer is yes, but this is a non-trivial task due to several challenges. (1) How can we model both watertight components (e.g., bodies, shoes, and accessories) and non-watertight, simulation-ready garments within a unified framework? While more expressive representations (e.g., G-Shell [42]) offer potential solutions, initializing and regularizing them to accurately represent open surfaces without multi-view image inputs remains an open problem. (2) Providing suitable constraints for optimizing decoupled bodies and outfits to achieve domain-specific realism is both key and challenging.
@@
To address these challenges, we introduce DreamBarbie [^1] Barbie® is a registered trademark of Mattel, Inc. This project is for academic research purposes only and is not affiliated with, endorsed by, or connected to Mattel, Inc., a novel text-driven framework for generating Barbie-style 3D avatars. As shown in Fig. LABEL:fig:teaser, avatars from DreamBarbie not only exhibit exquisite geometry and textures but also feature a variety of decoupled shoes, accessories, and simulation-ready garments that can be freely combined and edited. Furthermore, our framework supports expressive animation and compatibility with physical simulations, thereby satisfying all four aforementioned requirements simultaneously.
The core of the framework consists of two key components: (1) Unlike prior methods, we utilize expressive G-Shell [42] to unify watertight and non-watertight modeling. We propose an SDF-based boundary initialization that avoids manifold geodesics via a 3D proxy (Pie Mesh), ensuring stable gradients and $100\times$ acceleration. Combined with a hole-preserving loss, this enables clean open surface modeling without multi-view supervision. (2) Instead of relying on a single general model as in prior text-to-disentangled-avatar methods, we appropriately incorporate different expert diffusion models to guarantee the domain-specific fidelity. A series of effective regularization losses and strategies is also proposed to address geometric artifacts and texture conflicts when combining different expert models. Specifically, our method generates Barbie-like 3D digital humans in three stages: First, we generate a reasonable and realistic base body using human-specific diffusion models along with a proposed SMPLX-evolving prior loss. Second, we initialize apparel with the semantic-aligned body and optimize it using object-specific generative priors and several solid geometric losses. Finally, we jointly fine-tune the assembled avatar to enhance texture harmony and consistency. Through this pipeline, the generated 3D avatars are animatable and disentangled, featuring high-quality bodies, shoes, accessories, and simulation-ready garments, delivering a truly immersive “Barbie-style” digital experience.
@@
  • We introduce DreamBarbie, a novel text-driven framework for generating realistic and highly disentangled 3D avatars. The framework enables the decoupling of bodies, garments, shoes, and accessories, while enabling outfit transfer, editing, expressive animation, and simulation.
@@
  • Extensive experiments demonstrate that DreamBarbie outperforms existing methods in avatar and outfit generation, achieving superior results in geometry quality, texture detail, and alignment with text descriptions.
@@ II Related Work
Text to Holistic 3D Avatar Generation. The field of text-to-3D generation has witnessed significant progress [52, 7, 75, 65, 69, 38, 54, 62], largely driven by advances in text-to-image models [56, 58, 79]. Building upon the success in generating general 3D objects, various methods [19, 29, 23, 78, 3, 41, 76, 24, 39, 45, 4, 22, 73] have been proposed to generate complex 3D avatars. AvatarCLIP [19] pioneers zero-shot generation of 3D digital humans from text prompts by leveraging human priors and CLIP [55]. With the introduction of the Score Distillation Sampling (SDS) loss [52], several subsequent works [29, 23, 78, 3] combine it with parametric human models [43, 51, 2] to significantly improve the fidelity of generated avatars. To enhance animation expressiveness and simulation compatibility, approaches [41, 76, 24, 39, 45, 4] replace implicit NeRF [46, 67] with explicit representations such as 3DGS [27] and SMPL-X [51] for modeling 3D humans. However, purely implicit or explicit representations struggle to capture complex geometric details, limiting the quality of generated digital humans. Accordingly, HumanNorm [22], SeeAvatar [73], and AvatarStudio [80] propose using the hybrid representation DMTet [61], which combines the advantages of both implicit and explicit representations to achieve high-fidelity avatar generation. Nevertheless, these methods typically model the human body, garments, shoes, and accessories as a single model, sacrificing flexibility and controllability for downstream applications such as apparel composition and physical simulation.
@@
In contrast, avatars from DreamBarbie not only exhibit exquisite geometry and appearance but also wear multiple separable, realistic outfits, while supporting expressive animation and physical simulation. We summarize the main differences between DreamBarbie and existing methods in Table I.
@@ III-B Overview
Given a text prompt, DreamBarbie aims to create an animatable, disentangled 3D avatar dressed in simulation-ready garments, along with diverse shoes and accessories, resembling iconic Barbie dolls. Specifically, our framework consists of three stages: (1) Human Body Generation: This stage generates a reasonable and realistic basic human body by leveraging human-specific generative priors and a novel SMPLX-evolving prior loss (Sec. III-C and Fig. 2); (2) Apparel Generation: This stage models high-quality garments, shoes, and accessories piece by piece, utilizing object-specific diffusion models together with several initialization strategies and geometric losses (Sec. III-D and Fig. 3); and (3) Unified Texture Refinement: This stage enhances visual harmony and consistency by jointly fine-tuning the composed avatar (Sec. III-E and Fig. 3).
@@ III-C Human Body Generation
Since the human body is a closed surface, we utilize G-Shell ${\theta}_{h}$ in watertight mesh mode (i.e., ignoring mSDF and retaining only SDF) and a texture field ${\psi}_{h}$ to model the geometry and appearance of the human body, respectively. These components are optimized under human-specific generative priors and a novel SMPLX-evolving prior loss, as illustrated in Fig. 2.
Human Body Geometry Modeling. To enhance the robustness of 3D human generation, we first optimize the shape parameters $\beta$ of an SMPL-X model using SDS loss, aligning the coarse body shape with the input text description. The resulting initial mesh, denoted $M_{init}$, serves two key purposes: (1) initializing the human G-Shell $\theta_{h}$, and (2) providing rich, semantically aligned human priors to guide realistic outfit generation. As discussed in the introduction, using a general T2I model struggles to provide domain-specific constraints necessary for creating realistic human bodies or outfits. To address this limitation, we employ human-specific diffusion models from HumanNorm [22], which are fine-tuned with high-fidelity human data.
@@
To address this problem, we propose the SMPLX-evolving prior loss inspired by the evolving constraint introduced in SeeAvatar [73]. Specifically, we freeze all parameters of the SMPL-X model and enhance the initial mesh $M_{init}$ into an evolvable representation $\hat{M}_{init}$ by adding learnable vertex-wise offsets. These offsets are periodically fitted to the current body every $\delta$ iteration, allowing the optimized offsets to model complex body features. The fitting objective is as follows:
@@
where $\mathcal{L}_{chamf}$ is the Chamfer distance between $\hat{M}_{init}$ and the current body shape, $\mathcal{L}_{edge}$, $\mathcal{L}_{nor}$ and $\mathcal{L}_{lap}$ are the edge length regularization, normal consistency loss, and Laplacian smoothness, respectively. The advantages of this loss are twofold: (1) Enhanced Modeling Capabilities: Compared to the smooth $M_{init}$, our $\hat{M}_{init}$ excels at capturing detailed geometric features, such as muscle contours (see the zoomed area at the bottom of Fig. 2), and offers reliable yet diverse priors for subsequent geometry generation (Fig. 8-(a)). (2) Topology Preservation: Unlike the evolving templates with arbitrary topology in SeeAvatar [73], our $\hat{M}_{init}$ preserves the SMPL-X model’s topology and semantics during the evolution process. This makes it particularly suitable for apparel initialization, composition, animation, and simulation (Fig. 9).
@@ III-D Apparel Generation
Given an SMPLX-aligned basic body generated in the previous stage, we proceed to dress the Barbie-like avatar in various outfits. Similar to body modeling, we employ G-Shell ${\theta}_{a}$ for apparel geometry and a texture field ${\psi}_{a}$ for apparel texture. Closed shoes and accessories are represented using G-Shell in watertight mesh mode (i.e., ignoring mSDF and retaining only SDF), while open garments are modeled using G-Shell in non-watertight mesh mode (i.e., retaining both SDF and mSDF). As illustrated in Fig. 3, the apparel generation process begins by initializing apparel using predefined SMPL-X masks that cover over a dozen daily outfits. We then create high-quality apparel piece by piece using object-specific generative priors and well-designed geometric losses. Lastly, we fine-tune the assembled avatar to improve appearance harmony and consistency.
@@
Apparel Initialization. In the absence of multi-view supervision [42, 5], text-driven optimization of G-Shell representations heavily depends on effective initialization, especially for open surface modeling. Therefore, we carefully design two strategies for initializing open clothes and closed shoes/accessories, enabling accurate topology modeling for each category. (1) Closed Surface Initialization: We crop the open sub-mesh of the human body using the corresponding SMPL-X mask, expand it along vertex normals to create inner and outer layers, and sew them to form a double-layer closed template mesh $M_{temp}$. This mesh initializes the SDF $s_{\theta_{a}}(\cdot)$ of the apparel G-Shell ${\theta}_{a}$. Fig. 4-(a) shows this process. (2) Open Surface Initialization: As mSDF requires a closed manifold domain, we first construct a single-layer watertight template, $M_{temp}$, by hole-filling the cropped body mesh. This simplified topology serves as the canonical domain for the mSDF. The primary challenge lies in defining the open boundaries on $M_{temp}$. While a geodesic-based initialization (i.e., signed geodesic distance to boundaries) is intuitive, it is computationally prohibitive ($\sim$5 hours) and mathematically ill-posed for 3D implicit fields. Because geodesics are confined to the 2-manifold, they lack a stable gradient flow in 3D space, resulting in optimization artifacts (Fig. 4-(c)).
To address this, we reformulate boundary modeling as a field-based intersection problem. Specifically, we introduce the Pie Mesh ($M_{pie}$), a localized watertight proxy that encapsulates each hole’s boundary (see Supp. Mat. for fitting details). Mathematically, instead of manifold traversal, we use $M_{pie}$’s SDF values to initialize the mSDF $\hat{s}_{\theta_{a}}(\cdot)$. This ensures that points sampled inside $M_{pie}$ have negative mSDF values (indicating they lie inside the open surface), while those outside have positive values (indicating they lie outside the open surface). Consequently, the open surface is implicitly determined by the set subtraction between the $M_{temp}$ manifold and the $M_{pie}$ volume (Fig. 4-(b), black dotted box). Fig. 4-(b) illustrates the whole process. This strategy offers three advantages: (1) Efficiency: Replacing manifold-constrained path-finding with SDF queries achieves a $100\times$ speedup ($\sim$3 mins vs. $\sim$5 hours, see the Supp. Mat. for speed comparison details). (2) Stability: The SDF provides a globally continuous gradient field, ensuring stable supervision and eliminating floating artifacts or unsmooth boundaries. (3) Regularization: $M_{pie}$ serves as a geometric prior, providing explicit boundary constraints for subsequent optimization (see Eq. 14). Fig. 4-(c) compares these two initialization methods.
@@
where $y_{a}$ is the description of a single outfit, and $n^{a}$ and $d^{a}$ are the rendered normal and depth maps of the apparel in the undressed state. Besides, $\mathcal{L}^{{sdn}}_{SDS}$ is the vanilla Stable Diffusion (SD) SDS loss enforced on the rendered apparel normal maps. As shown in RichDreamer [54], naive SD helps object-specific diffusion models produce more stable results. In this way, two specific generative priors provide powerful and reasonable guidance to modeling high-fidelity outfits (Fig. 8-(b)).